A garden is in the from of a trapezium whose paralles are 40 m and 22 m and the perpendicular distance between them is 12 m. find the area of the garden ?

Grade:8

4 Answers

rathnakar
6 years ago
 
A garden is in the from of a trapezium whose paralles are 40 m and 22 m and the perpendicular distance between them is 12 m. find the area of the garden ?
rathnakar
6 years ago
A garden is in the from of a trapezium whose paralles are 40 m and 22 m and the perpendicular distance between them is 12 m. find the area of the garden ?
Nimit Jain
6 years ago
Since area of trapezium is given by the formula: 1/2*(sum of parallel sides)*hArea = 1/2*(40+22)(12) ((The perpendicular distance will be the height)) = 66*6 = 396 sq.m
Ft
6 years ago
area of trapezium = 1/2*(sum of parallel sides)*h so ,Area = 1/2*(40+22)(12) ((The perpendicular distance will be the height)) area of trapezium garden=1/2*62*12sq.m =372sq.m